Could you please help me solve this question? One day you sewed two regular book bags and four deluxe book bags and it took you a total of 12.4 hrs. The next day, you sewed three regular and two deluxe and it took a total of 9.4 hrs. How much time does one regular bag take, and how much time does one deluxe bag take? Thank you.
Answer by checkley75(3666)

2R+4D=12.4
3R+2D=9.4 NOW MULTIPLY THIS EQUATION BY -2 AND ADD THESE 2 EQUATIONS.
-6R-4D=-18.8
2R+4D=12.4
--------------------------------
-4R=-6.4
R=-6.4/-4
R=1.6 HOURS TO MAKE A REGULAR BAG.
2*1.6+4D=12.4
3.2+4D=12.4
4D=12.4-3.2
4D=9.2
D=9.2/4
2.3 HOURS FOR THE DELUX BAG.
PROOF
3*1.6+2*2.3=9.4
4.8+4.6=9.4
9.4=9.4